OpenStudy (anonymous):

Find the value of a: -3/7 = a/14 a = 6 a = -6 a = 1/6 a = -1/6

OpenStudy (anonymous):

@skittles_for_life6422

OpenStudy (freckles):

There is thing people like to call "cross multiply". It isn't really a mathematical term but you can do that.

OpenStudy (freckles):

it really is just multiplying the bottoms on both sides

OpenStudy (anonymous):

I dont get it

OpenStudy (freckles):

or if you want we can "reduce" -3/7 I put quotations around the reduce because we are actually going to be multiply top and bottom of -3/7 by an integer such that we can get blank/14

OpenStudy (freckles):

\[\frac{-3}{7}=\frac{?}{14}\] what number times 7 is 14?

OpenStudy (anonymous):

so it would be -1/6

OpenStudy (freckles):

7 times what =14

OpenStudy (anonymous):

2?

OpenStudy (freckles):

yes so if you multiply the bottom by 2 you multiply the top by 2

OpenStudy (freckles):

\[\frac{-3}{7} \cdot \frac{2}{2}=?\]

OpenStudy (anonymous):

-6/14?

OpenStudy (freckles):

yep \[\frac{-3}{7}=\frac{-3}{7} \cdot \frac{2}{2}=\frac{-3(2)}{7(2)}=\frac{-3(2)}{14}=\frac{-6}{14}\]

OpenStudy (freckles):

so in short you have \[\frac{-3}{7}=\frac{-6}{14}\]

OpenStudy (anonymous):

but -6/14 is not an option?

OpenStudy (freckles):

I know because that isn't the answer

OpenStudy (freckles):

they are looking for the numerator

OpenStudy (freckles):

\[\frac{-3}{7}=\frac{?}{14} \\ \text{ and you have } \\ \frac{-3}{7}=\frac{6}{14}\]

OpenStudy (freckles):

so the ? is definitely not the whole fraction

OpenStudy (freckles):

it is just the numerator of that one fraction

OpenStudy (freckles):

oops forgot my negative sign up there

OpenStudy (freckles):

\[\frac{-3}{7}=\frac{?}{14} \\ \text{ and you have } \\ \frac{-3}{7}=\frac{-6}{14}\]

OpenStudy (freckles):

so compare these what should the question mark be?

OpenStudy (anonymous):

OH!